# KNN-MMD: Cross Domain Wireless Sensing via Local Distribution Alignment

### v2

@@ Abstract
Wireless sensing has recently found widespread applications in diverse environments, including homes, offices, and public spaces. By analyzing patterns in channel state information (CSI), it is possible to infer human actions for tasks such as person identification, gesture recognition, and fall detection. However, CSI is highly sensitive to environmental changes, where even minor alterations can significantly distort the CSI patterns. This sensitivity often leads to performance degradation or outright failure when applying wireless sensing models trained in one environment to another. To address this challenge, Domain Alignment (DAL) has been widely adopted for cross-domain classification tasks, as it focuses on aligning the global distributions of the source and target domains in feature space. Despite its popularity, DAL often neglects inter-category relationships, which can lead to misalignment between categories across domains, even when global alignment is achieved. To overcome these limitations, we propose K-Nearest Neighbors Maximum Mean Discrepancy (KNN-MMD), a novel few-shot method for cross-domain wireless sensing. Our approach begins by constructing a “help set” using K-Nearest Neighbors (KNN) from the target domain, enabling local alignment between the source and target domains within each category using Maximum Mean Discrepancy (MMD). Additionally, we address a key instability issue commonly observed in cross-domain methods, where model performance fluctuates sharply between epochs. Further, most existing methods struggle to determine an optimal stopping point during training due to the absence of labeled data from the target domain. Our method resolves this by excluding the support set from the target domain during training and employing it as a validation set to determine the stopping criterion. We evaluate the effectiveness of the proposed method across several cross-domain Wi-Fi sensing tasks, including gesture recognition, person identification, fall detection, and action recognition, using both a public dataset and a self-collected dataset. In a one-shot scenario, our method achieves accuracy rates of 93.26%, 81.84%, 77.62%, and 75.30% for the respective tasks. To support future research, we will release our code and dataset to the public upon publication.
@@
Few-shot Learning, K-Nearest Neighbors, Maximum Mean Discrepancy, Cross-domain Wi-Fi Sensing, Channel Statement Information
@@ I Introduction
Coverage Privacy Cost
Camera Low (affected by occlusion/lighting) Low Moderate
Millimeter Wave Moderate Moderate High
Infrared Sensor Moderate High Moderate
Wi-Fi High (non-line-of-sight, all-day) High Low
Given the released tool that works with some specific types of commercial Network Interface Cards (NIC) [5], the Channel State Information (CSI) can be obtained, which describes how the wireless signals fade during the propagation from the sender to the receiver. By using this CSI data collection tool, a series of research employing commercial Wi-Fi NICs for different kinds of tasks. Moreover, a dedicated IEEE standard for Wi-Fi sensing named IEEE 802.11bf is approved [6] to enable faster development and practicalization of Wi-Fi sensing technologies. However, since the Orthogonal Frequency Division Multiplexing (OFDM) waveforms employed by Wi-Fi are designed for communication purpose without sensing optimization, it is thus challenging to apply physical models to infer and sense the ambient environment from the CSI measurements.
@@
As the most common method for cross-domain tasks, Domain Adaptation (DA) has been widely utilized in cross-domain Wi-Fi sensing. DA focuses on leveraging knowledge from a source domain to enhance model performance on a related but different target domain. Among these methods, Domain Alignment (DAL), which aims to map the source and target domains to the same distribution, is one of the most popular approaches and has been extensively applied in Wi-Fi sensing [11, 12, 13, 14, 15, 16, 17]. However, as shown in Fig. 1, the global alignment of traditional DAL methods can lead to incorrect alignments within categories, which can hinder the effectiveness of the classification boundary trained in the source domain when applied to the target domain. Another widely recognized issue in DA is low stability: the model’s performance on the testing set can be unstable during training (as referenced in [18] and Section IV-D), with accuracy varying sharply between successive epochs. This instability suggests that simply stopping the model training at a fixed epoch may result in poor performance. If we wish to implement an early stopping method, we require labeled validation data from the target domain. However, the only labeled data available from the target domain is the support set, which is often a subset of the training set and cannot be used as a validation set.
@@ II-A Wi-Fi Sensing Principle
In Wi-Fi sensing, CSI is one of the most commonly used features. The channel between the transmitter and receiver can be influenced when encountering dynamic or static objects. The RX will then receive the multipath superimposed signals from refraction, reflection, and scattering. By analyzing the received CSI, we can obtain information about the environment and behavior. The channel model can be mathematically represented as:
@@ II-B Domain Shift Issue in Wi-Fi Sensing
Figure 2: Average and Standard Deviation of CSI Amplitude (WiGesture Dataset) [3]: Each point represents a $1$s sample. The x-axis represents different people and action IDs, while different colors represent different actions.
To facilitate understanding, we first define the symbols and the cross-domain scenarios. In the following, the variables $x$ and $y$ refer to the input data and output label, respectively. $\theta$ represents the network parameters. $P_{s}$ and $P_{t}$ represent the classification probabilities of each category in the source domain and target domain, respectively. Furthermore, the focus of this paper is primarily on the few-shot cross-domain task. Most available data is in the source domain expect $n$ labeled samples for each category in the target domain. This scenario is often referred to as the n-shot problem. In addition, we also have some unlabeled data from the target domain. Our objective is to accurately classify these unlabeled data. For the sake of simplicity, we will refer to the labeled data from the source domain as the training set, and the labeled data from the target domain as the support set.
@@ II-C Existing Domain Adaptation Methods
DAL methods aim to map the source domain and target domain to a common feature distribution, where the data in the two domains can have the same distribution in the new space. However, most DAL methods, such as MMD [25], Kernel Mean Matching (KMM) [34], and Subspace Geodesic Flow (SGF) [35], rely on the assumption that the conditional probabilities in the source domain and target domain are the same, i.e., $P_{t}(y|x)=P_{s}(y|x)=P(y|x)$. To estimate $P(y|x)$, the network parameters $\theta$ need to be learned to achieve $P(y|x;\theta)=P(y|x)$. However, due to $P_{t}(x)\neq P_{s}(x)$, the $P(y|x;\theta)$ learned from the source domain cannot generalize to the target domain. Consequently, these traditional methods can be seen as finding a latent space where the input data from the source domain and target domain are mapped to the same distribution, i.e., $P_{t}(x)=P_{s}(x)$. This allows the model to learn the parameter $\theta$ more effectively and accurately. However, in many practical scenarios, the assumption $P_{t}(y|x)=P_{s}(y|x)$ is not met. This has been demonstrated in our experiments (Section IV-C) as well as our previous research [18], where traditional DAL methods even failed in cross-domain Wi-Fi sensing tasks.
@@ III-A Overview
As shown in Fig. 4, our method mainly has two steps: (1) Preliminary Classification and (2) Training Network with Local Alignment. In Step 1, we use KNN to preliminarily classify the testing set according to the support set. We then choose the samples with the highest classification confidence levels to construct a help set for Step 2. The help set will be used for the local alignment with the training set. In Step 2, we train a classification network with the MK-MMD loss function to align the source domain and the target domain within each category. Additionally, we use an early stop strategy according to the classification accuracy in the support set, as it does not participate in the training and has the same distribution as the testing set. By this way, our training process can stop at a point where the testing set has a relatively high classification accuracy.
@@ III-B Local Distribution Alignment
Let’s reconsider the cross-domain classification task. As proposed in [38], the assumption $P_{t}(y|x)=P_{s}(y|x)$ is not met in most practical scenarios. As a result, we consider the case when $P_{t}(y|x)\neq P_{s}(y|x)$ and aim to find a feature space where, after mapping, $P_{t}(y|x^{\prime})=P_{s}(y|x^{\prime})$. In this feature space, we can train a classifier network to realize the final classification.
We can use the conditional probability formula to structure the relationship between $P_{s}(y|x)$ and $P_{t}(y|x)$ as:
$P_{s}(y|x)$ $=P_{t}(y|x)\frac{P_{s}(x|y)P_{s}(y)}{P_{s}(x)}\frac{P_{t}(x)}{P_{t}(x|y)P_{t}(y)}$ (3)
$=P_{t}(y|x)\frac{P_{t}(x)}{P_{s}(x)}\frac{P_{s}(x|y)}{P_{t}(x|y)}\frac{P_{s}(y)}{P_{t}(y)}$
$=P_{t}(y|x)\frac{\sum_{y^{\prime}}P_{t}(y^{\prime})P_{t}(x|y^{\prime})}{\sum_{y^{\prime}}P_{s}(y^{\prime})P_{s}(x|y^{\prime})}\frac{P_{s}(x|y)}{P_{t}(x|y)}\frac{P_{s}(y)}{P_{t}(y)}\ ,$
where $y^{\prime}$ represents the index of the category in the mapped feature space. In most cases, we have $P_{s}(y)\approx P_{t}(y)$, indicating that the label distribution is similar between the source and target domains. If we also have $P_{t}(x|y)\approx P_{s}(x|y)$, then $P_{s}(y|x)\approx P_{t}(y|x)$.
@@
where $K$ is the set of kernel functions provided by the user, and $\beta$ is a set of weights that satisfy $\sum_{h=1}^{H}\beta_{h}=1$. During training, the MK-MMD loss function can be used in the embedding result of the source and target domains to encourage the feature extractor to map the source and target domain data to the same distribution in the intermediate feature space.
@@ III-C Model Structure
Our model architecture is illustrated in Fig. 6, which primarily consists of three components: a normalization layer, a feature extractor, and a final classifier. The normalization layer incorporates time-level standardization and batch normalization, helping to reduce the domain gap between different samples and facilitating the training process. Then we select the pre-trained ResNet18 [20] as the feature extractor because the Wi-Fi signal spectrum exhibits high similarity to images. Convolutional networks, which effectively extract local features, have demonstrated excellent performance in CSI classification [18, 1, 40]. Furthermore, inspired by transfer learning, the pre-trained parameters from image data provide a strong starting point for Wi-Fi sensing [41]. Finally, we employ a MLP to process the embeddings for classification. The detailed description is as follows:
First, we use a standard layer[3] as shown in Eq. 7:
$\text{Std}(X_{i}^{(j)})=\frac{X_{i}^{(j)}-E_{time}[X^{(j)}]}{\sqrt{D_{time}[X^{(j)}]}}\ ,$ (7)
where $X_{i}^{(j)}$ represents the $i$-th sample in the $j$-th subcarrier of the CSI sequence $X$, $E$ and $D$ denote the expectation and variance, respectively. The standardization operation is applied in the time dimension. In [3], it’s noted that the distribution of CSI would have a significant difference in different time ranges, which could make the network difficult to converge. By standardization, the data distribution in the time dimension can be unified to a similar distribution, which helps to solve this problem, to a certain extent.
Then, the data is input to a batch normalization layer that normalizes each element along the batch dimension, as shown in Eq. 8:
$\text{BN}(X)=\gamma\frac{X-E_{batch}[X]}{\sqrt{D_{batch}[X]}}+\beta\ ,$ (8)
where $\gamma$ and $\beta$ are learnable parameters. Batch normalization is a traditional method to deal with the covariate shift problem[42], which is similar to domain shift but with a smaller domain gap. As a result, we use it to deal with the domain shift problem.
Subsequently, the processed data is fed into Resnet to obtain the embedding, which serves as the feature space mapping mentioned in Section III-B. Finally, the MLP acts as a classifier, taking the embedding as input and producing the corresponding label as output. The network can be expressed as Eq. 9.
$E$ $=\text{Resnet}(\text{BN}(\text{Std}(X)))\ ,$ (9)
@@
where $E$ is the embedding result, and $\hat{Y}$ is the output.

### v3

@@ Abstract
Wireless sensing has recently found widespread applications in diverse environments, including homes, offices, and public spaces. By analyzing patterns in channel state information (CSI), it is possible to infer human actions for tasks such as person identification, gesture recognition, and fall detection. However, CSI is highly sensitive to environmental changes, where even minor alterations can significantly distort the CSI patterns. This sensitivity often leads to performance degradation or outright failure when applying wireless sensing models trained in one environment to another. To address this challenge, Domain Alignment Learning (DAL) has been widely adopted for cross-domain classification tasks, as it focuses on aligning the global distributions of the source and target domains in feature space. Despite its popularity, DAL often neglects inter-category relationships, which can lead to misalignment between categories across domains, even when global alignment is achieved. To overcome these limitations, we propose K-Nearest Neighbors Maximum Mean Discrepancy (KNN-MMD), a novel few-shot method for cross-domain wireless sensing. Our approach begins by constructing a “help set” using K-Nearest Neighbors (KNN) from the target domain, enabling local alignment between the source and target domains within each category using Maximum Mean Discrepancy (MMD). Additionally, we address a key instability issue commonly observed in cross-domain methods, where model performance fluctuates sharply between epochs. Further, most existing methods struggle to determine an optimal stopping point during training due to the absence of labeled data from the target domain. Our method resolves this by excluding the support set from the target domain during training and employing it as a validation set to determine the stopping criterion. We evaluate the effectiveness of the proposed method across several cross-domain Wi-Fi sensing tasks, including gesture recognition, person identification, fall detection, and action recognition, using both a public dataset and a self-collected dataset. In a one-shot scenario, our method achieves accuracy rates of 93.26%, 81.84%, 77.62%, and 75.30% for the respective tasks. The dataset and code are publicly available at https://github.com/RS2002/KNN-MMD.
@@
Few-shot Learning, K-Nearest Neighbors, Maximum Mean Discrepancy, Cross-domain Wi-Fi Sensing, Channel State Information
@@ I Introduction
Functionality Accuracy Coverage Privacy Cost
Camera Comprehensive Excellent Limited (affected by occlusion/lighting) Exposed Relatively Inexpensive
Millimeter Wave Fairly Comprehensive Quite Good Moderate Fairly Secure Expensive
Infrared Sensor Limited Excellent Moderate Secure Inexpensive
Wi-Fi Fairly Comprehensive Quite Good Good (non-line-of-sight, all-day) Secure Negligible
Given the released tool that works with some specific types of commercial Network Interface Cards (NIC) [5], the Channel State Information (CSI) can be obtained, which describes how the wireless signals fade during the propagation from the sender to the receiver. By using this CSI data collection tool, a series of research employing commercial Wi-Fi NICs for different kinds of tasks. Moreover, a dedicated IEEE standard for Wi-Fi sensing named IEEE 802.11bf is approved [6] to enable faster development and practicalization of Wi-Fi sensing technologies. However, since the Orthogonal Frequency Division Multiplexing (OFDM) waveforms employed by Wi-Fi are designed for communication purposes without sensing optimization, it is thus challenging to apply physical models to infer and sense the ambient environment from the CSI measurements.
@@
As the most common method for cross-domain tasks, Domain Adaptation (DA) has been widely utilized in cross-domain Wi-Fi sensing. DA focuses on leveraging knowledge from a source domain to enhance model performance on a related but different target domain. Among these methods, Domain Alignment Learning (DAL), which aims to map the source and target domains to the same distribution, is one of the most popular approaches and has been extensively applied in Wi-Fi sensing [11, 12, 13, 14, 15, 16, 17]. However, as shown in Fig. 1, the global alignment of traditional DAL methods can lead to incorrect alignments within categories, which can hinder the effectiveness of the classification boundary trained in the source domain when applied to the target domain. Another widely recognized issue in DA is low stability: the model’s performance on the testing set can be unstable during training (as referenced in [18] and Section IV-D), with accuracy varying sharply between successive epochs. This instability suggests that simply stopping the model training at a fixed epoch may result in poor performance. If we wish to implement an early stopping method, we require labeled validation data from the target domain. However, the only labeled data available from the target domain is the support set, which is often a subset of the training set and cannot be used as a validation set.
@@ II-A Wi-Fi Sensing Principle
In Wi-Fi sensing, CSI is one of the most commonly used features. The channel between the transmitter and receiver can be influenced when encountering dynamic or static objects. The receiver (RX) will then receive the multipath superimposed signals from refraction, reflection, and scattering. By analyzing the received CSI, we can obtain information about the environment and behavior. The channel model can be mathematically represented as:
@@ II-B Domain Shift Issue in Wi-Fi Sensing
Figure 2: Average and Standard Deviation of CSI Amplitude (WiGesture Dataset) [3]: Each point represents a 1 second sample. The x-axis represents different people and action IDs, while different colors represent different actions.
To facilitate understanding, we first define the symbols and the cross-domain scenarios. In the following, the variables $x$ and $y$ refer to the input data and output label, respectively. $\theta$ represents the network parameters. $P_{s}$ and $P_{t}$ represent the classification probabilities of each category in the source domain and target domain, respectively. Furthermore, the focus of this paper is primarily on the few-shot cross-domain task. Most available data is in the source domain except $n$ labeled samples for each category in the target domain. This scenario is often referred to as the n-shot problem. In addition, we also have some unlabeled data from the target domain. Our objective is to accurately classify these unlabeled data. For the sake of simplicity, we will refer to the labeled data from the source domain as the training set, and the labeled data from the target domain as the support set.
@@ II-C Existing Domain Adaptation Methods
DAL methods aim to map the source domain and target domain to a common feature distribution, allowing the data in both domains to share the same distribution in the new space. However, most DAL methods, such as MMD [25], Kernel Mean Matching (KMM) [34], and Subspace Geodesic Flow (SGF) [35], rely on the assumption that the conditional probabilities in the source and target domains are the same, i.e., $P_{t}(y|x)=P_{s}(y|x)=P(y|x)$, where $P_{s}$ and $P_{t}$ represent the probability distributions in the source and target domains, respectively, and $x$ and $y$ represent the data sample and ground truth category label. When training the network in the source domain, the network parameters $\theta$ need to be learned to achieve $P(y|x;\theta)\approx P(y|x)$ under the source domain $P_{s}(x)$. However, since $P_{t}(x)\neq P_{s}(x)$, the $P(y|x;\theta)$ learned from the source domain cannot generalize to the target domain $P_{t}(x)$. Consequently, these traditional methods can be viewed as attempting to find a latent space where the input data from the source and target domains are mapped to the same distribution, i.e., $P_{t}(\text{g}(x))\approx P_{s}(\text{g}(x))$, where $\text{g}(\cdot)$ is a mapping function. This allows the model to learn the parameters $\theta$ more effectively and accurately in the target domain. However, in many practical scenarios, the assumption $P_{t}(y|x)=P_{s}(y|x)$ is not satisfied. This has been demonstrated in our experiments (Section IV-C) as well as in our previous research [18], where traditional DAL methods have even failed in cross-domain Wi-Fi sensing tasks. A more detailed derivation of the formulas can be found in Section III-B.
@@
Remark 1.
Although some previous work has highlighted the disadvantages of conventional global alignment and proposed similar pseudo-label-based DAL methods, we want to emphasize the differences between our work and theirs.
First, we argue that some prior DAL methods still struggle in challenging scenarios. For instance, the method proposed by [36] cannot effectively address the scenario illustrated in Fig. 1, as mentioned in the previous section. Our proposed method, which directly aligns the source and target domains within each category, represents a simple yet effective approach to achieve local alignment.
The success of our method primarily relies on the quality of pseudo labels for the target domain. An inevitable issue in pseudo-label-based methods is the occurrence of incorrect classifications and noise in the data. Most previous work [37, 38, 39] first trains a network in the source domain and then uses it to classify the target domain, generating pseudo labels. Typically, these methods retain only the top classification samples from the target domain for subsequent DAL training, aiming to minimize noise and incorrect primary classifications as much as possible. While these methods have shown considerable success in fields like CV and electrocardiogram (ECG) analysis, the Wi-Fi CSI serves as a different modality, often exhibiting a completely different data distribution between source and target domains. This discrepancy can render the network trained on the source domain ineffective in the target domain. For example, using the WiGesture dataset, we train a ResNet18 in the source domain and test its performance in the target domain. The confusion matrix is shown in Fig. 4a. For category 2-4, most classification results are incorrect. Notably, in category 2, there are no correct generated labels, which hinders subsequent operations and leads to significantly low model performance. In Fig. 4b, we also display the accuracy across different categories under the top $p\%$ classification confidence level. For classes other than 0 and 4, regardless of how $p\%$ is selected, the accuracy remains low, resulting in poor quality of pseudo labels. In contrast, our method utilizes model-based KNN for pseudo label generation, which is less affected by cross-domain issues compared to neural networks. Additionally, we propose a confidence-level-based approach to mitigate the influence of noise and incorrect classifications, which has been demonstrated to be effective through sensitivity analysis.
Furthermore, our method supports an early stopping mechanism, which is not feasible for most DAL methods. This feature allows us to efficiently determine when to stop the model at the optimal point, significantly enhancing model stability.
(a) Confusion Matrix
(b) Accuracy under Top $p\%$ Confidence Level
Figure 4: Model Performance of Vanilla ResNet18 on WiGesture Dataset
@@ III-A Overview
As shown in Fig. 5, our method mainly has two steps: (1) Preliminary Classification and (2) Training Network with Local Alignment. In Step 1, we use KNN to preliminarily classify the testing set according to the support set. We then choose the samples with the highest classification confidence levels to construct a help set for Step 2. The help set will be used for the local alignment with the training set. In Step 2, we train a classification network with the Multiple Kernel Maximum Mean Discrepancy (MK-MMD) loss function to align the source domain and the target domain within each category. Additionally, we use an early stop strategy according to the classification accuracy in the support set, as it does not participate in the training and has the same distribution as the testing set. By this way, our training process can stop at a point where the testing set has a relatively high classification accuracy.
@@ III-B Local Distribution Alignment
Let’s reconsider the cross-domain classification task. During training, we have a labeled dataset from the source domain, denoted as $D_{s}=(X_{s},Y_{s})$, and an unlabeled dataset from the target domain, represented as $D_{t}=X_{t}$. Here, $X_{s}$ and $X_{t}$ are the sample data from the source and target domains, respectively, while $Y_{s}$ and $Y_{t}$ denote their category labels. When we train a network using $D_{s}$, we obtain the network parameters $\theta$ through Maximum Likelihood Estimation (MLE):
$\theta=\arg\max_{\theta}P(Y_{s}|\text{f}(X_{s};\theta))\ ,$ (3)
where $\text{f}(\cdot;\theta)$ represents the neural network. However, due to the significant domain gap between the source and target domains, the network cannot guarantee the probability of samples in the target domain, i.e., $P(Y_{t}|\text{f}(X_{t},\theta))$. In conventional DAL methods, a mapping function $\text{g}(\cdot)$ is introduced to align the data from the source and target domains to a similar distribution, such that $P(\text{g}(X_{s}))\approx P(\text{g}(X_{t}))$, which we refer to as global alignment.
However, we note that this approach does not ensure model performance in the target domain. Consequently, the network learns the parameters $\theta$ as follows:
$\theta=\arg\max_{\theta}P(Y_{s}|\text{f}(\text{g}(X_{s});\theta))\ .$ (4)
This can be interpreted as the network fitting the label distribution in the source domain:
$P(y|\text{f}(\text{g}(x);\theta))\approx P_{s}(y|\text{g}(x))\ ,$ (5)
where $x$ and $y$ are arbitrary data samples and their corresponding ground truth labels, respectively. The distributions $P_{s}$ and $P_{t}$ represent the probabilities in the source and target domains, respectively. For the network to function effectively in the target domain, we require that $P_{s}(y|\text{g}(x))\approx P_{t}(y|\text{g}(x))$. According to Bayes’ rule, we have:
$P_{s}(y|\text{g}(x))$ $=P_{t}(y|\text{g}(x))\frac{P_{s}(\text{g}(x)|y)P_{s}(y)}{P_{s}(\text{g}(x))}\frac{P_{t}(\text{g}(x))}{P_{t}(\text{g}(x)|y)P_{t}(y)}$ (6)
$=P_{t}(y|\text{g}(x))\frac{P_{t}(\text{g}(x))}{P_{s}(\text{g}(x))}\frac{P_{s}(\text{g}(x)|y)}{P_{t}(\text{g}(x)|y)}\frac{P_{s}(y)}{P_{t}(y)}$
$=P_{t}(y|\text{g}(x))\frac{\sum_{y^{\prime}\in\tilde{Y}}P_{t}(y^{\prime})P_{t}(\text{g}(x)|y^{\prime})}{\sum_{y^{\prime}\in\tilde{Y}}P_{s}(y^{\prime})P_{s}(\text{g}(x)|y^{\prime})}$
$\quad\cdot\frac{P_{s}(\text{g}(x)|y)}{P_{t}(\text{g}(x)|y)}\frac{P_{s}(y)}{P_{t}(y)}\ ,$
where $\tilde{Y}$ denotes the set of all categories. In global alignment, we only have $P_{s}(\text{g}(x))\approx P_{t}(\text{g}(x))$. However, this alone is insufficient to ensure $P_{s}(y|\text{g}(x))\approx P_{t}(y|\text{g}(x))$. In most cases, we find that $P_{s}(y)\approx P_{t}(y)$, indicating that the label distributions are similar between the source and target domains. If we also have $P_{t}(\text{g}(x)|y)\approx P_{s}(\text{g}(x)|y)$, then it follows that $P_{s}(y|\text{g}(x))\approx P_{t}(y|\text{g}(x))$. We refer to this sufficient condition as local alignment.
@@
where $K$ is the set of kernel functions provided by the user, $\beta$ is a set of weights that satisfy $\sum_{h=1}^{H}\beta_{h}=1$, and $H$ is the kernel amount. During training, the MK-MMD loss function can be used in the embedding result of the source and target domains to encourage the feature extractor to map the source and target domain data to the same distribution in the intermediate feature space.
@@ III-C Model Structure
Our model architecture is illustrated in Fig. 6, which primarily consists of three components: a batch normalization layer, a feature extractor, and a final classifier. The model takes the standard CSI sequence as input, where the signal is first standardized over the time dimension, a method shown to be effective for CSI sensing model training [3]. In the network, the batch normalization layer helps reduce the domain gap between different samples, facilitating the training process. We then select the pre-trained ResNet18 [20] as the feature extractor, as the Wi-Fi signal spectrum exhibits high similarity to images. Convolutional networks, which effectively extract local features, have demonstrated excellent performance in CSI classification [18, 1, 41]. Furthermore, inspired by transfer learning, the pre-trained parameters from image data provide a strong starting point for Wi-Fi sensing [42]. Finally, we employ a multilayer perceptron (MLP) to process the embeddings for classification. The detailed description is as follows:
After batch normalization, the processed data is fed into ResNet to obtain the embedding, which serves as the feature space mapping mentioned in Section III-B. Finally, the MLP acts as a classifier, taking the embedding as input and producing the corresponding label as output. The network can be expressed as Eq. 10.
$E$ $=\text{ResNet}(\text{BatchNorm}(X))\ ,$ (10)
@@
where $X$ is the input, $E$ is the embedding result, and $\hat{Y}$ is the output.